InTheSea:

What is the total power drawn by the circuit shown? 2.14 kW 2.14 W 22.0 kW 20.0 W

InTheSea:

@Shadow (110/20)+(110/8)+(110/500)=19.47 I took a wrong turn at Albuquerque

1 attachment
Shadow:

FUN

Shadow:

(110/20)+(110/8)+(110/500)=19.47 What is this

InTheSea:

110/ohms

Shadow:

You are dividing voltage, by the resistor. Why

Shadow:

For total current?

InTheSea:

Yea

Shadow:

\[P = IV\]

Shadow:

Your work is correct, that's the total current. Keep going.

InTheSea:

I don't know where to go from here

Shadow:

Power = Current times Voltage

InTheSea:

19.47*110=2141.7 watts=2.1417 kw so A?

Shadow:

Correct